OpenStudy (anonymous):

how do you solve a complex numbers questions where z conjugate+2z = -4-4i+10+3i, where you want to solve for z

OpenStudy (anonymous):

dY=dX + ZdX + XdZ-2*ZdZ. =(1+Z)dX + (X-2Z)dZ. 3.2) Y=40 + 3X + 4XZ. 2. dY=3dX + 4Z. 2. dX + 8XZdZ. = (3+4Z. 2 ). dX +8XZdZ. 3.3) For 3.1),

OpenStudy (anonymous):

what is dY?

OpenStudy (helder_edwin):

\[ \large \overline{z}+2z=-4-4i+10+3i=6-i \] let \(z=a+ib\) so \(\overline{z}=a-ib\) then \[ \large 6-i=\overline{z}+2z=(a-ib)+2(a+ib)=3a+ib \] so \(3a=6\) and \(b=-1\) and from this \(a=2\) and \(b=-1\) therefore \[ \large z=2-i \]

OpenStudy (anonymous):

THANK YOU SO MUCH